# Characterizations of graph classes via convex geometries: A survey

### v6

@@ 3.4 $m^{3}$-convexity
Let $P_{n}$ denote the induced path with $n$ vertices. An alternative definition of “simplicial” is: $v$ is a simplicial vertex iff it is not a midpoint of an induced $P_{3}$. In jamison-olariu , this concept is relaxed as follows: a vertex is semisimplicial if it is not an internal vertex of an induced $P_{4}$, and nonsemisimplicial otherwise. Clearly, a vertex $v$ is an extreme vertex of an $m^{3}$-convex set $S\subseteq V(G)$ if and only if $v$ is semisimplicial in $G[S]$.
A hole is an induced cycle with at least five vertices. The house is the graph with vertices $a,b,c,d,e$ and edges $ab,bc,cd,ad,ae,be$. The domino is the graph with vertices $a,b,c,d,e,f$ and edges $ab,bc,cd,ad,ce,ef,df$. Finally, the $A$ is the domino minus the edge $ef$. A graph $G$ is weakly polarizable olariu if $G$ contains no hole, house, domino, or $A$ as an induced subgraph. Weakly polarizable graphs are also called HHDA-free graphs. Using arguments similar to those presented in the preceding subsections, it can be shown if the $m^{3}$-convexity of $G$ is a convex geometry then $G$ is a weakly polarizable graph. Suppose the $m^{3}$-convexity of $G$ is a convex geometry and $G$ contains, for instance, a house $G^{\prime}$ as an induced subgraph. Then the only possible semisimplicial vertex in the subgraph of $G$ induced by $Y=H(V(G^{\prime}))$ is vertex $e$, i.e., ${\mathit{e}xt}(Y)\subseteq\{e\}$. This implies $H({\mathit{e}xt}(Y))\neq Y$, a contradiction. Analogously, $G$ cannot contain a hole, a domino, or an $A$ as an induced subgraph. Thus, $G$ is weakly polarizable. Conversely, if $G$ is weakly polarizable, it holds that every nonsemisimplicial vertex lies on an induced path of length at least three between semisimplicial vertices dragan-et-al . We then have the following theorem:
@@
dragan-et-al A graph $G$ is weakly polarizable if and only if the $m^{3}$-convexity of $G$ is a convex geometry.
@@ 3.5 Toll convexity
A walk $u_{0}u_{1}\ldots u_{k-1}u_{k}$ is a tolled walk if $u_{0}u_{i}\in E(G)$ implies $i=1$ and $u_{j}u_{k}\in E(G)$ implies $j=k-1$. In alcon , tolled walks were used to characterize interval graphs. A graph $G$ is an interval graph if its vertices can be associated with intervals on the real line such that two vertices are adjacent if and only if the associated intervals intersect.
@@
alcon-et-al If $v$ is an extreme vertex of a toll convex set $S$ of a graph $G$ then $v$ is simplicial in $G[S]$.
@@
alcon-et-al A vertex $v$ of an interval graph $G$ is an extreme vertex of a toll convex set $S\subseteq V(G)$ if and only if $v$ is an end simplicial vertex of $G[S]$.
@@
lek-bol A graph $G$ is an interval graph if and only if $G$ is chordal and contains no asteroidal triple.
Suppose that the toll convexity of $G$ is a convex geometry. Let us show first that $G$ is chordal. By the assumption, $V(G)$ (and every of its toll convex subsets) is the convex hull of its extreme vertices. Let $v$ be an extreme vertex of $V(G)$. By definition of extreme vertex, $V(G)\setminus\{v\}$ is toll convex in $G$. Thus, the toll convexity of $G-v$ is a convex geometry (note that any subset of $V(G)\setminus\{v\}$ is toll convex in $G-v$ iff it is toll convex in $G$). Using induction, $G-v$ is chordal. Since by Lemma 8 $v$ is a simplicial vertex of $G$, it follows that $G$ is chordal as well (recall that a graph is chordal iff there is an ordering $v_{1},\ldots,v_{n}$ of its vertices such that $v_{i}$ is simplicial in $G[\{v_{i},v_{i+1},\ldots,v_{n}\}]$, $1\leq i\leq n$ golumbic80 ).
@@
The precedent paragraphs tell us that if the toll convexity of $G$ is a convex geometry then, by Theorem 10, $G$ is an interval graph. To prove the converse, we use the following useful lemma:
@@
alcon-et-al Let $G$ be an interval graph. Then every vertex that is not an end simplicial vertex lies in a tolled walk between two end simplicial vertices of $G$.
@@
alcon-et-al A graph $G$ is an interval graph if and only if the toll convexity of $G$ is a convex geometry.
@@ 3.6 Weakly toll convexity
A walk $u_{0}u_{1}\ldots u_{k-1}u_{k}$ is a weakly toll walk if $u_{0}u_{i}\in E(G)$ implies $u_{i}=u_{1}$ and $u_{j}u_{k}\in E(G)$ implies $u_{j}=u_{k-1}$. The concept of weakly toll walk is a relaxation of the concept of tolled walk (in the sense that every tolled walk is a weakly toll walk). In the previous subsection, tolled walks have been used to characterize interval graphs via convex geometries. Similarly, weakly toll walks are used as a tool to characterize proper interval graphs. A proper interval graph is an interval graph that admits an interval model in which no interval properly contains another, or, equivalently, an interval model in which all the intervals have the same length. Roberts roberts proved that proper interval graphs are exactly the interval graphs containing no $K_{1,3}$ as an induced subgraph. The graph $K_{1,3}$ consists of four vertices $a,b,c,d$ and three edges $ab,ac,ad$.
Let ${\mathscr{P}}^{w}$ be the path system that maps a graph $G$ to its weakly toll walks. The convexity defined over such walks is called the weakly toll convexity of $G$, and the associated convex sets are called weakly toll convex sets.
Lemmas 8 and 9 have similar versions for weakly toll convex sets:
@@
gutierrez-tondato If $v$ is an extreme vertex of a weakly toll convex set $S$ of a graph $G$ then $v$ is simplicial in $G[S]$.
@@
gutierrez-tondato A vertex $v$ of a proper interval graph $G$ is an extreme vertex of a weakly toll convex set $S\subseteq V(G)$ if and only if $v$ is an end simplicial vertex of $G[S]$.
Using arguments similar to those used in the previous section, one can prove that if the weakly toll convexity of $G$ is a convex geometry then $G$ is chordal and cannot contain asteroidal triples and induced subgraphs isomorphic to $K_{1,3}$. Hence, using the characterization of proper interval graphs in roberts , $G$ is a proper interval graph. Conversely, assume that $G$ is a proper interval graph. Then, it is an interval graph. This means that, by Lemma 11, every vertex of $G$ that is not an end simplicial vertex lies in a tolled walk between two end simplicial vertices. Since every tolled walk is a weakly toll walk and, by Lemma 14, end simplicial vertices of a proper interval graph are extreme vertices, we have:
@@
gutierrez-tondato Let $G$ be a proper interval graph. Then every vertex that is not an end simplicial vertex lies in a weakly toll walk between two end simplicial vertices of $G$.
@@
gutierrez-tondato A graph $G$ is a proper interval graph if and only if the weakly toll convexity of $G$ is a convex geometry.
@@ 3.7 $l^{k}$-convexity
Let ${\mathscr{P}}^{\leq k}$ be the path system that maps a graph $G$ to its induced paths of length at most $k$. The convexity defined over such paths is called the $l^{k}$-convexity of $G$, and the associated convex sets are called $l^{k}$-convex sets. The $l^{2}$-convexity is also called $P_{3}^{*}$-convexity araujo .
Suppose that the $l^{2}$-convexity of $G$ is a convex geometry. It is easy to see that the extreme vertices of the $l^{2}$-convex set $V(G)$ are the simplicial vertices of $G$. Thus, if $G$ contains an induced cycle $C$ with at least four vertices then $H(V(C))$ does not have simplicial (extreme) vertices, a contradiction. This means that $G$ is chordal. Now, we prove that $G$ is also a cograph (a graph that contains no $P_{4}$ as an induced subgraph – see corneil-et-al for details). Let $P=x_{0},\ldots,x_{p}$ be a maximum induced path of $G$, between two simplicial vertices of $G$. Such a path exists because $G$ is chordal dirac . Assume $p\geq 3$. The set $H(V(P))$ is an $l^{2}$-convex set of $G$ whose extreme vertices are $x_{0}$ and $x_{p}$. If there exists $z\in H(V(P))\setminus V(P)$ that lies in an induced path of length two between $x_{0}$ and $x_{p}$, $z$ must be adjacent to every $x_{i}$; otherwise, $P$ plus the edges $x_{0}z$ and $zx_{p}$ forms an induced cycle of $G$ with at least four vertices. Since $N[x_{0}]$ and $N[x_{p}]$ are cliques in $H(V(P))$,
@@
araujo-sampaio ; gpt A graph $G$ is a chordal cograph if and only if the $l^{2}$-convexity of $G$ is a convex geometry.
We now analyze graphs with $l^{3}$-convexities that are convex geometries. A vertex $u$ is a universal vertex in a graph $G$ if $u$ is adjacent to every other vertex of $G$. An $n$-gem is a graph $G_{n}$ such that: (i) $V(G_{n})=\{x_{0},\ldots,x_{n},u_{n}\}$ ($n\geq 4$); (ii) $x_{0},\ldots,x_{n}$ is an induced path; (iii) $u_{n}$ is a universal vertex. We say that $G_{n}$ is solved if there exists in $G$ a $P_{4}$ connecting $x_{0}$ and $x_{n}$ that avoids $u_{n}$. In gpt , graphs with $l^{3}$-convexities that are convex geometries are characterized as follows:
@@
gpt Let $G$ be a graph. The $l^{3}$-convexity of $G$ is a convex geometry if and only if the following conditions hold:

### v7

@@ 3.4 $m^{3}$-convexity
An alternative definition of “simplicial” is: $v$ is a simplicial vertex iff it is not a midpoint of an induced $P_{3}$. In [31], this concept is relaxed as follows: a vertex is semisimplicial if it is not an internal vertex of an induced $P_{4}$, and nonsemisimplicial otherwise. Clearly, a vertex $v$ is an extreme vertex of an $m^{3}$-convex set $S\subseteq V(G)$ if and only if $v$ is semisimplicial in $G[S]$.
A graph $G$ is weak bipolarizable [36] if $G$ contains no hole, house, domino, or $A$ as an induced subgraph. See Figure 3. Weak bipolarizable graphs are also called HHDA-free graphs.
Figure 3: From left to right: hole, house, domino, $A$.
Using arguments similar to those presented in the preceding subsections, it can be shown if the $m^{3}$-convexity of $G$ is a convex geometry, then $G$ is a weak bipolarizable graph. Suppose the $m^{3}$-convexity of $G$ is a convex geometry and $G$ contains, for instance, a house $G^{\prime}$ as an induced subgraph. Then, the only possible semisimplicial vertex in the subgraph of $G$ induced by $Y=H(V(G^{\prime}))$ is vertex $e$, i.e., ${\mathit{e}xt}(Y)\subseteq\{e\}$. This implies $H({\mathit{e}xt}(Y))\neq Y$, a contradiction. Analogously, $G$ cannot contain a hole, a domino, or an $A$ as an induced subgraph. Thus, $G$ is weak bipolarizable. Conversely, if $G$ is weak bipolarizable, it holds that every nonsemisimplicial vertex lies on an induced path of length at least three between semisimplicial vertices [19]. We then have the following theorem:
@@
[19] A graph $G$ is weak bipolarizable if and only if the $m^{3}$-convexity of $G$ is a convex geometry.
3.5 Steiner convexity
Let $S\subseteq V(G)$. A Steiner tree for $S$ is a connected subgraph $T$ of $G$ such that $S\subseteq V(T)$ and $|E(T)|$ is minimum. The Steiner interval $I(S)$ is the set formed by all vertices of $G$ that lie in some Steiner tree for $S$. We say that $S$ is Steiner convex if, for every $S^{\prime}\subseteq S$, it holds that $I(S^{\prime})\subseteq S$. The Steiner convexity of $G$ is the family formed by the Steiner convex sets of $G$. Important references on the Steiner convexity are [9, 16].
In [9], the authors prove that: (a) the extreme vertices of a Steiner convex set $S$ of $G$ are the simplicial vertices of $G[S]$; (b) in a (house,hole,domino)-free graph $G$, the Steiner convexity and the geodesic convexity coincide. Such facts are the main ingredients to prove the following characterization:
Theorem 8.
[9] A graph $G$ is Ptolemaic if and only if the Steiner convexity of $G$ is a convex geometry.
Proof.
Suppose that $G$ is Ptolemaic, i.e., $G$ is chordal and gem-free. Then $G$ is clearly (house,hole,domino)-free, so the Steiner convexity of $G$ coincides with the geodesic convexity of $G$, which is a convex geometry by Theorem 2.
Conversely, suppose that the Steiner convexity of $G$ is a convex geometry. We proceed with arguments that we have used before. If $G$ contains an induced gem $G^{\prime}$, then, clearly, the extreme vertices of $Y=H(V(G^{\prime}))$ are in $\{a,d\}$ (see Figure 2). However, $H(\{a,d\})\neq Y$. Thus, $G$ is gem-free. Likewise, if $G$ contains an induced cycle $C$ with at least four vertices then $W=H(V(C))$ is Steiner convex and contains no simplicial vertices, i.e., ${\mathit{e}xt}(W)=\emptyset$. Thus, $G$ is chordal. ∎
In [10], an alternative way of defining Steiner convexity is proposed, by looking at subsets with a prescribed size. We say that $S$ is $k$-Steiner convex or $g_{k}$-convex if, for every $S^{\prime}\subseteq S$ with $|S^{\prime}|=k$, it holds that $I(S^{\prime})\subseteq S$. The convexity associated with the $g_{k}$-convex sets of $G$ is called the $k$-Steiner convexity of $G$. Note that the $2$-Steiner convexity of $G$ is precisely the geodesic convexity of $G$.
Let us study the case $k=3$. In [10], the extreme vertices of $g_{3}$-convex sets are characterized: a vertex $x$ is an extreme vertex of a $g_{3}$-convex set $S$ of $G$ if and only if $x$ is not the central vertex of an induced claw, paw, or $P_{4}$ contained in $G[S]$. See Figure 4.
[…]
Figure 4: From left to right: claw, paw, $P_{4}$. Central vertices are indicated by arrows.
Suppose that the $3$-Steiner convexity of $G$ is a convex geometry. Suppose also that $G$ contains an induced graph $P$ isomorphic to a $P_{4}=abcd$, where $a$ and $d$ have degree one in $P$. Observe that $a$ and $d$ are the only extreme vertices of $P$. Thus, ${\mathit{e}xt}(V(P))$ is a set $C\subseteq\{a,d\}$. However, $H(C)=C\neq H(\{a,b,c,d\})$, that is, $H(\{a,b,c,d\})$ is a 3-Steiner convex set that is not the convex hull of its extreme points. This is a contradiction. Hence, $G$ is $P_{4}$-free.
Now consider the four graphs $G_{1},G_{2},G_{3},G_{4}$ depicted in Figure 5. Note that ${\mathit{e}xt}(G_{i})=\emptyset,1\leq i\leq 4$. For instance, it can be checked by inspection that every vertex in $G_{1}$ is a central vertex of a claw. A similar behavior occurs for $G_{2}$, $G_{3}$, and $G_{4}$. This implies that $Y=H(G_{i})$ is a 3-Steiner convex set and ${\mathit{e}xt}(Y)=\emptyset$, a contradiction. Hence, $G$ is $(G_{1},G_{2},G_{3},G_{4})$-free.
On the other hand, in [10] the authors show that if $G$ is $(P_{4},G_{1},G_{2},G_{3},G_{4})$-free then the $3$-Steiner convexity of $G$ is a convex geometry; this proof is a bit laborious, and so it is omitted here.
Theorem 9.
[10] The $3$-Steiner convexity of $G$ is a convex geometry if and only if $G$ is $(P_{4},G_{1},G_{2},G_{3},G_{4})$-free.
[…]
Figure 5: Graphs $G_{1},G_{2},G_{3},G_{4}$.
Let =${\cal R}=\{k_{1},k_{2},\ldots,k_{t}\}$ be a set of positive integers with $k_{1}\leq\cdots\leq k_{t}$ and $k_{1}\in\{2,3\}$. Say that a set $S\subseteq V(G)$ is ${\cal R}$-Steiner convex if $S$ is $k_{i}$-Steiner convex for $1\leq i\leq t$. The convexity associated with ${\cal R}$-Steiner convex sets is the ${\cal R}$-Steiner convexity. Interestingly, the class of graphs for which the ${\cal R}$-Steiner convexity is a convex geometry is exactly the same as the class for which the $k_{1}$-Steiner convexity is a convex geometry:
Theorem 10.
[10] The following statements are true:
  1. 1.
    If $k=2$, then the ${\cal R}$-Steiner convexity is a convex geometry if and only if $G$ is Ptolemaic.
  2. 2.
    If $k=3$, then the ${\cal R}$-Steiner convexity is a convex geometry if and only if $G$ is $(P_{4},G_{1},G_{2},G_{3},G_{4})$-free.
To conclude this section, we briefly discuss a further convexity notion, also presented in [10], that combines the $3$-Steiner convexity and the so-called $g^{3}$-convexity. A set $S$ of vertices in a graph $G$ is $g^{3}$-convex if, for every pair $u,v\in S$ such that $u$ and $v$ are at distance at least $3$ in $G$, $I[u,v]\subseteq S$. (Here, the interval is taken over shortest paths.) A set $S$ is $g^{3}_{3}$-convex if it is both $g_{3}$- and $g^{3}$-convex. The convexity associated with $g^{3}_{3}$-convex sets is the $g^{3}_{3}$-Steiner convexity.
Theorem 11.
[…]
Figure 6: Graphs $T_{1}$ and $T_{2}$.
@@ 3.6 Toll convexity
A walk $u_{0}u_{1}\ldots u_{k-1}u_{k}$ is a tolled walk if $u_{0}u_{i}\in E(G)$ implies $i=1$ and $u_{j}u_{k}\in E(G)$ implies $j=k-1$. In [1], tolled walks were used to characterize interval graphs. A graph $G$ is an interval graph if its vertices can be associated with intervals on the real line such that two vertices are adjacent if and only if the associated intervals intersect. The key references on the toll convexity are [2, 15].
@@
[2] If $v$ is an extreme vertex of a toll convex set $S$ of a graph $G$, then $v$ is simplicial in $G[S]$.
@@
[2] A vertex $v$ of an interval graph $G$ is an extreme vertex of a toll convex set $S\subseteq V(G)$ if and only if $v$ is an end simplicial vertex of $G[S]$.
@@
[33] A graph $G$ is an interval graph if and only if $G$ is chordal and contains no asteroidal triple.
Suppose that the toll convexity of $G$ is a convex geometry. Let us show first that $G$ is chordal. By the assumption, $V(G)$ (and every of its toll convex subsets) is the convex hull of its extreme vertices. Let $v$ be an extreme vertex of $V(G)$. By definition of extreme vertex, $V(G)\setminus\{v\}$ is toll convex in $G$. Thus, the toll convexity of $G-v$ is a convex geometry (note that any subset of $V(G)\setminus\{v\}$ is toll convex in $G-v$ iff it is toll convex in $G$). Using induction, $G-v$ is chordal. Since by Lemma 12 $v$ is a simplicial vertex of $G$, it follows that $G$ is chordal as well (recall that a graph is chordal iff there is an ordering $v_{1},\ldots,v_{n}$ of its vertices such that $v_{i}$ is simplicial in $G[\{v_{i},v_{i+1},\ldots,v_{n}\}]$, $1\leq i\leq n$ [26]).
@@
The precedent paragraphs tell us that if the toll convexity of $G$ is a convex geometry, then, by Theorem 14, $G$ is an interval graph. To prove the converse, we use the following useful lemma:
@@
[2] Let $G$ be an interval graph. Then, every vertex that is not an end simplicial vertex lies in a tolled walk between two end simplicial vertices of $G$.
@@
[2] A graph $G$ is an interval graph if and only if the toll convexity of $G$ is a convex geometry.
@@ 3.7 Weakly toll convexity
A walk $u_{0}u_{1}\ldots u_{k-1}u_{k}$ is a weakly toll walk if $u_{0}u_{i}\in E(G)$ implies $u_{i}=u_{1}$ and $u_{j}u_{k}\in E(G)$ implies $u_{j}=u_{k-1}$. The concept of weakly toll walk is a relaxation of the concept of tolled walk (in the sense that every tolled walk is a weakly toll walk). In the previous subsection, tolled walks have been used to characterize interval graphs via convex geometries. Similarly, weakly toll walks are used as a tool to characterize proper interval graphs. A proper interval graph is an interval graph that admits an interval model in which no interval properly contains another, or, equivalently, an interval model in which all the intervals have the same length. Roberts [40] proved that proper interval graphs are exactly the interval graphs containing no $K_{1,3}$ as an induced subgraph. The graph $K_{1,3}$ consists of four vertices $a,b,c,d$ and three edges $ab,ac,ad$.
Let ${\mathscr{P}}^{w}$ be the path system that maps a graph $G$ to its weakly toll walks. The convexity defined over such walks is called the weakly toll convexity of $G$, and the associated convex sets are called weakly toll convex sets. The basic references on the weakly toll convexity are [28, 29].
Lemmas 12 and 13 have similar versions for weakly toll convex sets:
@@
[28] If $v$ is an extreme vertex of a weakly toll convex set $S$ of a graph $G$, then $v$ is simplicial in $G[S]$.
@@
[28] A vertex $v$ of a proper interval graph $G$ is an extreme vertex of a weakly toll convex set $S\subseteq V(G)$ if and only if $v$ is an end simplicial vertex of $G[S]$.
Using arguments similar to those used in the previous section, one can prove that if the weakly toll convexity of $G$ is a convex geometry, then $G$ is chordal and cannot contain asteroidal triples and induced subgraphs isomorphic to $K_{1,3}$. Hence, using the characterization of proper interval graphs in [40], $G$ is a proper interval graph. Conversely, assume that $G$ is a proper interval graph. Then, it is an interval graph. This means that, by Lemma 15, every vertex of $G$ that is not an end simplicial vertex lies in a tolled walk between two end simplicial vertices. Since every tolled walk is a weakly toll walk and, by Lemma 18, end simplicial vertices of a proper interval graph are extreme vertices, we have:
@@
[28] Let $G$ be a proper interval graph. Then, every vertex that is not an end simplicial vertex lies in a weakly toll walk between two end simplicial vertices of $G$.
@@
[28] A graph $G$ is a proper interval graph if and only if the weakly toll convexity of $G$ is a convex geometry.
@@ 3.8 $l^{k}$-convexity
Let ${\mathscr{P}}^{\leq k}$ be the path system that maps a graph $G$ to its induced paths of length at most $k$. The convexity defined over such paths is called the $l^{k}$-convexity of $G$, and the associated convex sets are called $l^{k}$-convex sets. The $l^{2}$-convexity is also called $P_{3}^{*}$-convexity [6].
Suppose that the $l^{2}$-convexity of $G$ is a convex geometry. It is easy to see that the extreme vertices of the $l^{2}$-convex set $V(G)$ are the simplicial vertices of $G$. Thus, if $G$ contains an induced cycle $C$ with at least four vertices, then $H(V(C))$ does not have simplicial (extreme) vertices, a contradiction. This means that $G$ is chordal. Now, we prove that $G$ is also a cograph (a graph that contains no $P_{4}$ as an induced subgraph – see [13] for details). Let $P=x_{0},\ldots,x_{p}$ be a maximum induced path of $G$, between two simplicial vertices of $G$. Such a path exists because $G$ is chordal [14]. Assume $p\geq 3$. The set $H(V(P))$ is an $l^{2}$-convex set of $G$ whose extreme vertices are $x_{0}$ and $x_{p}$. If there exists $z\in H(V(P))\setminus V(P)$ that lies in an induced path of length two between $x_{0}$ and $x_{p}$, $z$ must be adjacent to every $x_{i}$; otherwise, $P$ plus the edges $x_{0}z$ and $zx_{p}$ forms an induced cycle of $G$ with at least four vertices. Since $N[x_{0}]$ and $N[x_{p}]$ are cliques in $H(V(P))$,
@@
[7, 27] A graph $G$ is a chordal cograph if and only if the $l^{2}$-convexity of $G$ is a convex geometry.
We now analyze graphs with $l^{3}$-convexities that are convex geometries. A vertex $u$ is a universal vertex in a graph $G$ if $u$ is adjacent to every other vertex of $G$. An $n$-gem is a graph $G_{n}$ such that: (i) $V(G_{n})=\{x_{0},\ldots,x_{n},u_{n}\}$ ($n\geq 4$); (ii) $x_{0},\ldots,x_{n}$ is an induced path; (iii) $u_{n}$ is a universal vertex. See Figure 7.
$x_{0}$$x_{1}$$x_{n}$$u_{n}$
Figure 7: An $n$-gem.
Suppose that $G$ contains an $n$-gem $G_{n}$ as an induced subgraph. We say that $G_{n}$ is solved if there exists in $G$ a $P_{4}$ connecting $x_{0}$ and $x_{n}$ that avoids $u_{n}$. In [27], graphs with $l^{3}$-convexities that are convex geometries are characterized as follows:
@@
[27] Let $G$ be a graph. The $l^{3}$-convexity of $G$ is a convex geometry if and only if the following conditions hold: